# Intrinsic Dimension Correlation: uncovering nonlinear connections in multimodal representations

### v1

@@ Abstract
footnotetext:
Correspondence to: lorenzo.basile@phd.units.it.
Code is available at https://github.com/lorenzobasile/IDCorrelation
@@ 1 Introduction
Figure 1: Example usage of $I_{d}$or: we consider a $3$D dataset in which the points lie on the surface of a cylinder, hence whose intrinsic dimension ($I_{d}$) is $2$. We want to assess the correlation between the $2$D set of coordinates $xy$ (which also has $I_{d}=2$, as shown in the top-right panel) and the $1$D set $z$. Intuitively, as $x$ and $z$ describe a circle, it is evident that knowledge of $z$ (encoded by color) is very informative in determining the $x$ coordinate (e.g. a yellow point is sure to be found in the central region of the $x$ axis), but not in determining the $y$: hence, the correlation coefficient is $0.5$. Conversely, when estimating the correlation between $xz$ (whose $I_{d}$ is $1$, bottom-right panel) and $y$, we can see that having access to $y$ (which is now represented by color) does not give any information on the value of $x$ nor $z$, hence the correlation is 00.
@@
Modern machine learning models have the remarkable ability to extract subtle patterns from complex datasets and use them to perform a wide variety of tasks in an astonishingly accurate way. However, to date, we still lack a complete and accurate understanding of their inner workings, especially in the case of deep neural networks. An active field of research in the interpretability of neural networks is focused on characterizing and quantifying the similarity between different models. To this aim, many works [1, 2, 3] evaluate the statistical correlation between the latent representations produced by the models. This quantification is key because it allows, for example, to disentangle or tie together different aspects of data representations, allowing a better interpretation of how the model makes its decisions. Moreover, assessing the similarity between representations of different models is particularly useful to determine whether the latent spaces are compatible, meaning that information extracted by one model can be successfully transferred to others.
@@
This point is crucial particularly when data points are represented by multiple interrelated modalities, such as visual and textual. In recent times, it has been shown [4, 5] that it is possible to build multimodal vision-language models starting from pre-trained unimodal encoders, in an effort to match the outstanding performance of state-of-the-art vision-language models such as CLIP [6]. These findings indicate that modern deep models can produce compatible representations when evaluated on aligned text-image data. However, we find that standard latent similarity metrics such as CKA [2] and Distance Correlation [7, 8] find very frail traces of correlation between paired multimodal representations. We hypothesize that this is due to the strongly nonlinear nature of these correlations, making them difficult to analyze with standard methods.
@@
Our work tackles this challenge by introducing a novel metric, dubbed Intrinsic Dimension Correlation ($I_{d}$or) that leverages the concept of intrinsic dimension ($I_{d}$), i.e. the minimum number of variables required to describe the data, to quantify the correlation between high-dimensional data manifolds. Intuitively, the metric is based on the concept that if two data representations are correlated, the intrinsic dimension of a dataset created by concatenating the features of these representations is reduced, because the information in one representation can describe some aspects of the other. An example is provided in Fig. 1. Computational experiments indicate that our method is effective at detecting nonlinear relationships where traditional methods often fail.
@@
  • We propose a novel perspective that links the concepts of correlation and intrinsic dimension. The key concept is that being the $I_{d}$ a proxy for the information content of a dataset, the $I_{d}$ of a dataset built by merging two datasets will reflect the joint information.
@@ 2.1 Correlation in latent representations
Understanding whether different neural networks can learn to process data in similar ways is a crucial point when trying to make sense of their results. Despite the inherent difficulty in defining what it means for two neural networks to be similar (or, at least, to behave similarly), research in this field has made significant progress in the last few years. A well-established approach to test if two neural networks are similar is that of measuring the statistical correlation between the data representations (or embeddings) learned by them. Recent proposals in this direction include Singular Value Canonical Correlation Analysis (SVCCA) [1], Projection Weighted CCA (PWCCA) [9], Centered Kernel Alignment (CKA) [2], Distance Correlation [7, 8] and Aligned Cosine Similarity [10], among others. For a more complete summary of current approaches to neural network similarity measurement, we defer the reader to [11].
These techniques have been widely employed to gain a deeper understanding of various aspects of the way neural models process information: for instance to quantify how different vision architectures encode spatial information [12, 3] or the relation between learning disentangled features and adversarial robustness [8]. Taking a slightly different approach, [13] provides an in-depth analysis of the sensitivity of CKA to transformations that occur frequently in neural latent spaces, showcasing the importance of gathering results from a broader range of similarity metrics to obtain reliable information.
@@ 2.2 Multimodal latent space alignment
An example that highlights the importance of assessing similarity, quantified via correlation measures, between neural representations is further demonstrated by the recent empirical findings related to the so-called latent communication. This concept, introduced by [14], builds on the idea that it is possible to transfer knowledge between latent spaces, even when they are produced by different models and on different data modalities, provided that some semantic alignment between the data exists (for example, images and their textual descriptions). The feasibility of this knowledge transfer was shown in [14] through the introduction of relative representations, where each point of the original representation is mapped according to its distance from a set of fixed anchor points. Using this alternative representation of data, the authors show that it is possible to stitch [15] together encoders and decoders coming from different models, with little to no additional training.
Furthermore, numerous recent studies have demonstrated that large state-of-the-art visual and textual encoders can produce transferable representations when evaluated on aligned data (i.e. the same data or data that share some semantics, such as image-caption pairs). Indeed, a simple linear transformation is usually enough to map one latent space into another [5, 16, 17, 18], at least in terms of performance on a specific downstream task, e.g. classification. It is worth noting that, to perform the alignment of the data, one assumes prior knowledge about the semantic correlation between the data representations (in order to define the anchor points). Hence, while these findings suggest a remarkable similarity between compatible latent spaces, the problem of detecting these correlations without relying on any downstream evaluation is still an open problem. Indeed, our investigation into the connection between aligned textual and visual embeddings reveals a very weak correlation using existing methods, calling for the development of methods that allow the identification of nonlinear correlations in high-dimensional spaces.
@@ 2.3 Intrinsic dimension
Intrinsic dimension estimation is closely related to the field of dimensionality reduction since it gives a hint about which should be the dimension of the projection space to avoid information loss. Thus, one possible way of estimating the $I_{d}$ is to find a meaningful projection into the lowest dimensional space possible. A classical method for doing that is Principal Component Analysis [19], but it has the drawback that, strictly speaking, it is only correct if the data lie in a hyperplane, since it performs a linear transformation. Therefore, the development of methods for overcoming such a limitation is an active research field, resulting in techniques like Multidimensional Scaling [20], Isomap [21], t-distributed stochastic neighbor embedding (t-SNE) [22] or Uniform Manifold Approximation and Projection (UMAP) [23], to mention some. Other methods can estimate the $I_{d}$ of a dataset even in the case in which projecting in the lower dimensional space is not possible (for example, due to topological constraints). Typically, these approaches infer the $I_{d}$ from the properties of the Nearest Neighbors’ distances. While a full review of these methods is out of the scope of this work (the interested reader is referred to [24]), it is worth mentioning the Maximum Likelihood approach [25], the Dimensionality from Angle and Norm Concentration (DANCo) approach [26] or the TwoNN [27]. The last is the one employed in this work since it is particularly fast and behaves well even in the case of datasets with a high non-uniformity on the density of points.
More recently, several studies have estimated the intrinsic dimension of neural networks’ representations, demonstrating that $I_{d}$ is a valuable tool for understanding the geometry of the latent manifolds produced by deep models. This concept was initially explored in [28], where the authors estimated $I_{d}$ across different layers of CNNs, gaining insights into the sequential information flow within these models. Later, [29] and [30] analyzed the representations of transformer models, across different domains, while [31] studied the internal $I_{d}$ of generative diffusion models. In a slightly different direction, [32] unveiled a connection between generalization and the $I_{d}$ of the hidden representations, while [33] studied the relation between $I_{d}$ and curvature in latent manifolds.
@@ 3 Correlation through Intrinsic Dimension
The intrinsic dimension of a dataset is closely linked to the correlations among the various features that define the data points. These correlations determine the regions in which the data points can exist, thereby shaping the underlying manifold. Let us consider the simplest example: a two-dimensional dataset. If the two variables are uncorrelated, their linear correlation coefficient ($R^{2}$) approaches zero while, if one feature is a linear function of the other, $R^{2}$ becomes equal to one. The two scenarios differ by the $I_{d}$ of the data manifold: the first case corresponds to a plane ($I_{d}=2$), while the second corresponds to a line ($I_{d}=1$). If we examine a slightly more complex case, the advantage of exploiting the $I_{d}$ for correlation becomes evident. Let us consider a spiral-shaped 2D dataset: it has $R^{2}\approx 0$ due to the nonlinear nature of the correlation between the two variables, while the behavior of the $I_{d}$ is identical to the one observed on the linearly correlated dataset, as reported in detail in section 3.1.
To leverage these interesting properties for correlation quantification, we compute the $I_{d}$s of the two representations that we want to check for correlation ($id_{1}$ and $id_{2}$) and the $I_{d}$ of the dataset build by merging the two representations ($id_{C}$) and compute the correlation index $I_{d}$or as:
$I_{d}Cor=\frac{id_{1}+id_{2}-id_{C}}{\texttt{max}(id_{1},id_{2})}$ (1)
The intuition is that, if the two representations were independent, the $I_{d}$ of the merged dataset would be the sum of the two $I_{d}s$. Instead, if they were correlated, $id_{C}$ would be less then their sum since there would be some degree of redundancy between the coordinates of both representations.
A more formal view comes from information theory. Mutual information is a fundamental metric that quantifies the relationship between simultaneously sampled random variables. However, its computation from data samples is challenging due to the curse of dimensionality. One potential solution is the Sliced Mutual Information method [34], which projects high-dimensional random variables to scalars and averages over random projections. In our approach, presented in detail in Algorithm 1, we utilize the properties of intrinsic dimension to avoid the need for projection, thus examining the data in its entirety. Indeed, the intrinsic dimension can serve as a proxy for the information content of a dataset [35, 36]. Thus, $id_{1}$ and $id_{2}$ will be the information content of each of the datasets independently considered, while $id_{C}$ will be the information content of the joint dataset. Therefore, the numerator in the $I_{d}$or equation will correspond to the Mutual information between the datasets (applying the decomposition formula). Summarizing, $I_{d}$or is analogous to the Normalized Mutual Information (NMI) with max normalization [37], but it uses intrinsic dimension as a proxy for entropy. In this way, the correlation we adopt corresponds to the fraction of information contained in the more informative representation that is shared by the other dataset. Translated in terms of manifold learning, this represents the fraction of independent variables in the higher-dimensional dataset (in terms of $I_{d}$) that can be recovered from the lower-dimensional dataset.
To illustrate this, let us examine Fig. 1. In this synthetic 3D dataset, the data points lie on a cylinder-shaped manifold, where coordinates $x$ and $z$ are related through a nonlinear relationship, while $y$ is fully independent of the other two. This is evident when computing the $I_{d}$ values: the $I_{d}$ of the full dataset is $2$, as is the $I_{d}$ for the datasets with $xy$ or $yz$ variables, whereas the $I_{d}$ for the $xz$ dataset is $1$. Therefore, using the $I_{d}$or index to assess the correlation between $xy$ and $z$ yields a value of $0.5$, indicating that one of the two variables describing the $xy$ dataset can be recovered from $z$. Conversely, for the correlation between $xz$ and $y$, $I_{d}Cor=0$, confirming the statistical independence of $y$.
@@
Nevertheless, estimating $I_{d}$ can be prone to errors. To mitigate this, we developed a method to assign a $p$-value to the observed correlation. This method employs a permutation test [38] on $id_{C}$. Specifically, we calculate $I_{d}$ for several independent samples of a combined dataset, created by merging the two original datasets and randomizing the pairings to disrupt any existing correlations. This process allows us to determine the probability of accepting the hypothesis that the two datasets are uncorrelated.
Input : Two standardized datasets $\mathcal{D}_{1}\in\mathbb{R}^{N\times d_{1}}$ and $\mathcal{D}_{2}\in\mathbb{R}^{N\times d_{2}}$, an $I_{d}$ estimator id, and number of shuffles $S$.
Output : Correlation result $\varrho$ and $p$-value $p$
$id_{1}\leftarrow\texttt{id}(\mathcal{D}_{1})$;
$id_{2}\leftarrow\texttt{id}(\mathcal{D}_{2})$;
$id_{C}\leftarrow\texttt{id}(\texttt{concatenate}(\mathcal{D}_{1},\mathcal{D}_{2}))$;
$\varrho\leftarrow\frac{id_{1}+id_{2}-id_{C}}{\texttt{max}(id_{1},id_{2})}$;
$C\leftarrow 0$;
for $i\in\{1,...,S\}$ do
       $id_{s}[i]\leftarrow\texttt{id}(\texttt{concatenate}(\mathcal{D}_{1},\texttt{shuffle}(\mathcal{D}_{2})))$;
      if $id_{s}[i]\leq id_{C}$ then
             $C\leftarrow C+1$;
       end if
end for
$p\leftarrow\frac{C+1}{S+1}$;
return $\varrho$, $p$
Algorithm 1 $I_{d}$or

### v2

@@ 1 Introduction
Figure 1: Example usage of $I_{d}$or: we consider a $3$D dataset in which the points lie on the surface of a cylinder, hence whose intrinsic dimension ($I_{d}$) is $2$. We want to assess the correlation between the $2$D set of coordinates $xy$ (which also has $I_{d}=2$, as shown in the top-right panel) and the $1$D set $z$. Intuitively, as $x$ and $z$ describe a circle, it is evident that knowledge of $z$ (encoded by color) is very informative in determining the $x$ coordinate (e.g., a yellow point is sure to be found in the central region of the $x$ axis), but not in determining the $y$: hence, the correlation coefficient is $0.5$, according to equation 3. Conversely, when estimating the correlation between $xz$ (whose $I_{d}$ is $1$, bottom-right panel) and $y$, we can see that having access to $y$ (which is now represented by color) does not give any information on the value of $x$ nor $z$, hence the correlation is 00.
@@
Modern machine learning models have the remarkable ability to extract subtle patterns from complex datasets and use them to perform a wide variety of tasks in an astonishingly accurate way. However, to date, we still lack a complete and accurate understanding of their inner workings, especially in the case of deep neural networks. An active field of research in the interpretability of neural networks is focused on characterizing and quantifying the similarity between different models. To this aim, many works (Raghu et al., 2017; Kornblith et al., 2019; Nguyen et al., 2021) evaluate the statistical correlation between the latent representations produced by the models. This quantification is key because it allows, for example, to disentangle or tie together different aspects of data representations, allowing a better interpretation of how the model makes its decisions. Moreover, assessing the similarity between representations of different models is particularly useful to determine whether the latent spaces are compatible, meaning that information extracted by one model can be successfully transferred to others.
@@
This point is crucial particularly when data points are represented by multiple interrelated modalities, such as visual and textual. In recent times, it has been shown (Norelli et al., 2023; Moayeri et al., 2023) that it is possible to build multimodal vision-language models starting from pre-trained unimodal encoders, in an effort to match the outstanding performance of state-of-the-art vision-language models such as CLIP (Radford et al., 2021). These findings indicate that modern deep models can produce compatible representations when evaluated on aligned text-image data, hinting at a strong functional similarity between those. However, we find that standard latent similarity metrics such as CKA (Kornblith et al., 2019) and Distance Correlation (Székely et al., 2007; Zhen et al., 2022) find a very low structural correlation between paired multimodal representations. We hypothesize that this is due to the strongly nonlinear nature of these correlations, making them difficult to analyze with standard methods.
@@
Our work tackles this challenge by introducing a novel metric, dubbed Intrinsic Dimension Correlation ($I_{d}$or) that leverages the concept of intrinsic dimension ($I_{d}$), i.e., the minimum number of variables required to describe the data, to quantify the mutual information between high-dimensional data manifolds. Intuitively, the metric is based on the concept that if two data representations are correlated, the intrinsic dimension of a dataset created by concatenating the features of these representations is reduced, because the information in one representation can describe some aspects of the other. An example is provided in Fig. 1. It is worth noting that, to compute the correlation between two representations, a simpler approach based on the difference between the embedding dimension and the intrinsic dimension of the concatenated dataset would not suffice, since this would measure the amount of correlated features, regardless of whether these correlations are intra or inter representation. Computational experiments indicate that our method is effective at detecting nonlinear relationships where traditional methods often fail.
@@
  • We propose a novel perspective that links the concepts of correlation and intrinsic dimension. The key idea is that, since the $I_{d}$ serves as a proxy for the information content of a dataset, the $I_{d}$ of a dataset created by merging two datasets will represent their joint information.
@@ 2.1 Correlation in latent representations
Understanding whether different neural networks can learn to process data in similar ways is a crucial point when trying to make sense of their results. Despite the inherent difficulty in defining what it means for two neural networks to be similar (or, at least, to behave similarly), research in this field has made significant progress in the last few years. A well-established approach to test if two neural networks are similar is that of measuring the statistical correlation between the data representations (or embeddings) learned by them. Recent proposals in this direction include Singular Value Canonical Correlation Analysis (SVCCA) (Raghu et al., 2017), Projection Weighted CCA (PWCCA) (Morcos et al., 2018), Centered Kernel Alignment (CKA) (Kornblith et al., 2019), Distance Correlation (dCor) (Székely et al., 2007; Zhen et al., 2022), Aligned Cosine Similarity (Hamilton et al., 2016) and Representation Topology Divergence (RTD) (Barannikov et al., 2022), among others. For a more complete summary of current approaches to neural network similarity measurement, we defer the reader to Klabunde et al. (2023).
These techniques have been widely employed to gain a deeper understanding of various aspects of the way neural models process information: for instance to quantify how different vision architectures encode spatial information (Raghu et al., 2021; Nguyen et al., 2021) or the relation between learning disentangled features and adversarial robustness (Zhen et al., 2022). Taking a slightly different approach, Davari et al. (2023) provides an in-depth analysis of the sensitivity of CKA to transformations that occur frequently in neural latent spaces, showcasing the importance of gathering results from a broader range of similarity metrics to obtain reliable information.
@@ 2.2 Multimodal latent space alignment
An example that highlights the importance of assessing similarity, quantified via correlation measures, between neural representations is further demonstrated by the recent empirical findings related to the so-called latent communication. This concept, introduced by Moschella et al. (2023), builds on the idea that it is possible to transfer knowledge between latent spaces, even when they are produced by different models and on different data modalities, provided that some semantic alignment between the data exists (for example, images and their textual descriptions). The feasibility of this knowledge transfer was shown in Moschella et al. (2023) through the introduction of relative representations, where each point of the original representation is mapped according to its distance from a set of fixed anchor points. Using this alternative representation of data, the authors show that it is possible to stitch (Lenc & Vedaldi, 2015) together encoders and decoders coming from different models, with little to no additional training.
Furthermore, numerous recent studies have demonstrated that large state-of-the-art visual and textual encoders can produce transferable representations when evaluated on aligned data (i.e., the same data or data that share some semantics, such as image-caption pairs). Indeed, a simple linear transformation is usually enough to map one latent space into another (Moayeri et al., 2023; Merullo et al., 2023; Maiorca et al., 2023; Lähner & Moeller, 2024), at least in terms of performance on a specific downstream task, e.g., classification. It is worth noting that, to perform the alignment of the data, one assumes prior knowledge about the semantic correlation between the data representations (in order to define the anchor points). Hence, while these findings suggest a remarkable similarity between compatible latent spaces, the problem of detecting these correlations without relying on any downstream evaluation is still an open problem. Indeed, our investigation into the connection between aligned textual and visual embeddings reveals a very weak correlation using existing methods, calling for the development of methods that allow the identification of nonlinear correlations in high-dimensional spaces.
@@ 2.3 Intrinsic dimension
A possible way of estimating the $I_{d}$ is to find a meaningful projection (i.e., with minimal information loss) into the lowest dimensional space possible. A classical method for doing that is Principal Component Analysis (Wold et al., 1987), but it has the drawback that the intrinsic dimension it estimates is only correct if the manifold in which the data lie is a hyperplane. Therefore, the development of methods that can estimate the $I_{d}$ in nonlinear manifolds is an active research field. Typically, these approaches infer the $I_{d}$ from the properties of distances to the Nearest Neighbors. While a full review of these methods is out of the scope of this work (the interested reader is referred to Campadelli et al. (2015)), it is worth mentioning the Maximum Likelihood estimator (MLE) (Levina & Bickel, 2004), the Dimensionality from Angle and Norm Concentration (DANCo) approach (Ceruti et al., 2014) or the TwoNN (Facco et al., 2017). The last is the one employed in this work since it is particularly fast and behaves well even in the case of datasets with a high non-uniformity on the density of points. A brief description of TwoNN is provided in Appendix A.1.
More recently, several studies have estimated the intrinsic dimension of neural representations, demonstrating that $I_{d}$ is a valuable tool for understanding the geometry of the latent manifolds produced by deep models. This concept was initially explored in Ansuini et al. (2019), where the authors estimated $I_{d}$ across different layers of CNNs, gaining insights into the sequential information flow within these models. Later, Valeriani et al. (2023) and Cheng et al. (2023) analyzed the representations of transformer models, across different domains, while Kvinge et al. (2023) studied the internal $I_{d}$ of generative diffusion models. In a slightly different direction, Brown et al. (2022) unveiled a connection between generalization and the $I_{d}$ of the hidden representations, while Kaufman & Azencot (2023) studied the relation between $I_{d}$ and curvature in latent manifolds.
@@ 3 Correlation through Intrinsic Dimension
The intrinsic dimension of a dataset is closely linked to the correlations among the various features that define the data points. These correlations determine the regions in which the data points can exist, thereby shaping the underlying manifold. Let us consider the simplest example: a two-dimensional dataset. If the two variables are uncorrelated, their linear correlation coefficient ($R^{2}$) approaches zero while, if one feature is a linear function of the other, $R^{2}$ becomes equal to one. The two scenarios differ by the $I_{d}$ of the data manifold: the first case corresponds to a plane ($I_{d}=2$), while the second corresponds to a line ($I_{d}=1$). If we examine a slightly more complex case, the advantage of exploiting the $I_{d}$ for correlation becomes evident. Let us consider a spiral-shaped dataset embedded in two dimensions: it has $R^{2}\approx 0$ due to the nonlinear nature of the correlation between the two variables, while the behavior of the $I_{d}$ is identical to the one observed on the linearly correlated dataset, as reported in detail in section 4.1.
A formal framework that connects correlation and intrinsic dimension comes from information theory (Romano et al., 2016). In this field, mutual information is the fundamental metric that quantifies the relationship between simultaneously sampled random variables. Mutual information makes a natural candidate to serve as a measure of correlation between data, and Horibe (1985) defined a correlation coefficient between two random variables $X$ and $Y$ (known as Normalized Mutual Information - NMI) as:
$\varrho=\frac{I(X,Y)}{\max(H(X),H(Y))}$ (1)
where $I(\cdot,\cdot)$ is the mutual information and $H(\cdot)$ is the entropy.
In high dimensional datasets, computing the mutual information is challenging due to the curse of dimensionality, so we introduce the intrinsic dimension as a proxy. Mutual information can be expressed in terms of entropies as:
$I(X,Y)=H(X)+H(Y)-H(X,Y)$ (2)
Recent research (Bailey et al., 2022; Ghosh & Motani, 2023) has shown that intrinsic dimension is a metric that can satisfy the most desirable properties of entropy, while being easy and fast to compute on large-scale data (in Appendix A.2, we provide a test case where this relationship is exact). We exploit this relation to define our correlation coefficient $I_{d}$or, based on the NMI formulation, but using intrinsic dimension as a replacement for entropy. Given two standardized datasets (of possibly different dimensionality) $X\in\mathbb{R}^{n,d_{1}}$ and $Y\in\mathbb{R}^{n,d_{2}}$, we define their $I_{d}$or as:
$I_{d}\text{Cor}(X,Y)=\frac{(I_{d}(X)+I_{d}(Y)-I_{d}(X\oplus Y))}{\max(I_{d}(X),I_{d}(Y))}$ (3)
where $\oplus$ denotes row-wise concatenation. Although various other versions of NMI have been proposed over time (for a complete survey on variations of NMI we defer the reader to Vinh et al. (2009)) we adopt the $\max$ normalization since it allows an interpretation as the fraction of features of the most informative dataset that can be predicted using the less informative one (see Fig. 1 for an intuitive explanation of this concept).
@@
In practical applications, intrinsic dimension is calculated using estimators, which can be prone to errors. To mitigate this, we assign a $p$-value to the observed correlation, employing a permutation test (Davison & Hinkley, 1997) on $I_{d}(X\oplus Y)$. Specifically, we estimate the $I_{d}$ of several independent samples of the joint dataset, created by concatenating the two original datasets and randomizing the pairings to disrupt any existing correlations. This process allows us to determine a $p$-value that represents the probability of the hypothesis of the two datasets being uncorrelated as $p=\frac{L+1}{S+1}$, where $L$ is the number of estimates lower than the original joint $I_{d}$ and $S$ is the total number of permuted samples considered.
4 Results
In this section, we begin by assessing our proposed $I_{d}$or measure in simple, controlled settings using synthetic data. Then, we move to applications to latent representations generated by different neural networks on various datasets. We begin with a straightforward example that underscores the difficulty faced by standard methods in identifying nonlinear relationships, then progress to more extensive applications involving visual representations and multimodal text-image representations. For some experiments, we do not report the correlation $p$-value returned by our method. In such cases, the $p$-value is always the lowest possible according to the permutation test outlined in section 3 ($\frac{1}{S+1}$, where $S$ is the number of permutations, $100$ in most of our experiments).